Shift report 20030521 08:00-16:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Wed May 21 2003 - 16:15:21 EDT

  • Next message: Stephen J. Sanders: "brhijing"
    ------------------------------------------------------------
    Date: Wed May 21 16:15:20 2003, Shift: 08:00-16:00
    Supervisor: kh, with: 
    Log book#14 pages 115 - 115
    
    Shift Summary:
    ------------------------------------------------------------
    The pp2pp run was in progress today.  We turned on D5
    and the tracking detectors.  Took several min bias runs.
    FFS is at 8 deg (moved out of the way for pp2pp run); BFS is at 4 deg, so we took 0 field run for FFS.  D5 is set at 350B
    
    All in all, it is fairly laid back.
    
    ------------------------------------------------------------
    


    This archive was generated by hypermail 2.1.5 : Wed May 21 2003 - 16:19:18 EDT